OpenStudy (anonymous):

x-3/2x-(x-3)=x-2/(2x+3)-(x-3)

OpenStudy (anonymous):

Do you really want the answer to this problem, I mean do you need to show you work and all that good stuff

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

ok, I hope you have a printer for his one......

OpenStudy (anonymous):

x-(3)/(2)*x-(x-3)=x-(2)/(2x+3)-(x-3) Multiply -1 by each term inside the parentheses. x-(3)/(2)*x-(x-3)=x-(2)/(2x+3)-x+3 Since x and -x are like terms, add -x to x to get 0. x-(3)/(2)*x-(x-3)=0+3-(2)/(2x+3) Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. x-(3)/(2)*x-(x-3)=3-(2)/(2x+3) Multiply each term by a factor of 1 that will equate all the denominators. In this case, all terms need a denominator of (2x+3). x-(3)/(2)*x-(x-3)=3*(2x+3)/(2x+3)-(2)/(2x+3) Multiply the expression by a factor of 1 to create the least common denominator (LCD) of (2x+3). x-(3)/(2)*x-(x-3)=(3(2x+3))/(2x+3)-(2)/(2x+3) Multiply 3 by each term inside the parentheses. x-(3)/(2)*x-(x-3)=(6x+9)/(2x+3)-(2)/(2x+3) The numerators of expressions that have equal denominators can be combined. In this case, ((6x+9))/((2x+3)) and -(2)/((2x+3)) have the same denominator of (2x+3), so the numerators can be combined. x-(3)/(2)*x-(x-3)=((6x+9)-2)/(2x+3) Simplify the numerator of the expression. x-(3)/(2)*x-(x-3)=(6x+9-2)/(2x+3) Subtract 2 from 9 to get 7. x-(3)/(2)*x-(x-3)=(6x+7)/(2x+3) Multiply -1 by each term inside the parentheses. x-(3)/(2)*x-x+3=(6x+7)/(2x+3) Multiply -(3)/(2) by x to get -(3x)/(2). x-(3x)/(2)-x+3=(6x+7)/(2x+3) Since x and -x are like terms, add -x to x to get 0. 0-(3x)/(2)+3=(6x+7)/(2x+3) Combine all similar terms in the polynomial x-(3x)/(2)-x+3. -(3x)/(2)+3=(6x+7)/(2x+3) Find the LCD (least common denominator) of -(3x)/(2)+3+((6x+7))/((2x+3)). Least common denominator: 2(2x+3) Multiply each term in the equation by 2(2x+3) in order to remove all the denominators from the equation. -(3x)/(2)*2(2x+3)+3*2(2x+3)=(6x+7)/(2x+3)*2(2x+3) Simplify the left-hand side of the equation by canceling the common factors. -6x^(2)+3x+18=(6x+7)/(2x+3)*2(2x+3) Simplify the right-hand side of the equation by simplifying each term. -6x^(2)+3x+18=12x+14 Since 12x contains the variable to solve for, move it to the left-hand side of the equation by subtracting 12x from both sides. -6x^(2)+3x+18-12x=14 Since 3x and -12x are like terms, add -12x to 3x to get -9x. -6x^(2)-9x+18=14 To set the left-hand side of the equation equal to 0, move all the expressions to the left-hand side. -6x^(2)-9x+4=0 Multiply each term in the equation by -1. -6x^(2)*-1-9x*-1+4*-1=0*-1 Simplify the left-hand side of the equation by multiplying out all the terms. 6x^(2)+9x-4=0*-1 Multiply 0 by -1 to get 0. 6x^(2)+9x-4=0 Use the quadratic formula to find the solutions. In this case, the values are a=6, b=9, and c=-4. x=(-b\~(b^(2)-4ac))/(2a) where ax^(2)+bx+c=0 Use the standard form of the equation to find a, b, and c for this quadratic. a=6, b=9, and c=-4 Substitute in the values of a=6, b=9, and c=-4. x=(-9\~((9)^(2)-4(6)(-4)))/(2(6)) Simplify the section inside the radical. x=(-9\~(177))/(2(6)) Simplify the denominator of the quadratic formula. x=(-9\~(177))/(12) First, solve the + portion of \. x=(-9+~(177))/(12) Next, solve the - portion of \. x=(-9-~(177))/(12) The final answer is the combination of both solutions. x=(-9+~(177))/(12),(-9-~(177))/(12)_x<Z>APPR<z>0.359,-1.859

OpenStudy (anonymous):

i dont understand this

OpenStudy (anonymous):

hang on you just want the answer right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

can you minimize the steps???????????

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

X=(-9+√177)/12,(-9-√177)/12 or X≈0.359,-1.859 It will depend on the answer your teacher is looking for Has to how you will answer the equation.

OpenStudy (anonymous):

I hope this helps